OpenStudy (anonymous):

A 52.0-ml volume of 0.35M CH3COOH (Ka= 1.8*10^-5) is titrated with 0.40M NaOH . Calculate the pH after the addition of 21.0ml of NaOH.

OpenStudy (anonymous):

meq of CH3COOH=0.35*52=18.2. meq of NaOH= 21*0.4=8.4 meq of CH3COONa=8.4 meq of CH3COOH=9.8 By henderson's equation: pH=pKa + log salt/acid pH=5-log1.8+log 84/98=4.677